OpenStudy (anonymous):

The ratio of the heights of two similar rectangular prisms is 2:3. What is the ratio of their lateral areas? 8:27 2:5 4:9 1:0.33

jimthompson5910 (jim_thompson5910):

If the ratio of two sides is x:y then the ratio of the surface areas is x^2:y^2

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

what ^ means

jimthompson5910 (jim_thompson5910):

x^2 means \(\Large x^2\)

jimthompson5910 (jim_thompson5910):

the ^ tells the next number to go up and become an exponent

OpenStudy (anonymous):

oh ok lol

jimthompson5910 (jim_thompson5910):

sry I should have been more clear, let me know what you get for the answer

OpenStudy (anonymous):

so the correct answer is b

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

2^2 = ???

OpenStudy (anonymous):

oh c

jimthompson5910 (jim_thompson5910):

yes, 2^2 = 4 and 3^2 = 9 so if the sides are in ratio of 2:3, then the surface areas are in ratio of 4:9

OpenStudy (anonymous):

thanks for the medal and hellp

jimthompson5910 (jim_thompson5910):

you're welcome, you earned it by getting the correct answer and you hopefully see how to get it?